# Dissecting Buchdahl’s limit: A surgeon’s guide to compact objects

### v1

@@ III.1 Bilayered stars with positive core densities
Figure 2: Three dimensional representation of the surface in the $\{R_{i},\rho_{o},M\}$ space corresponding to solutions with an infinite central pressure and $\rho_{i}>0$. Every point on this surface lies above the standard Buchdahl limit $92M/R=8/$ 9, which is reached in the $\rho_{o}/\rho_{c}\to 1,~{}R_{i}/R\to 1$ limits. Lighter colours denote larger values of $2M/R$, and the black contour lines correspond to the $2M/R=\{0.92,0.94,0.96\}$ planes. This surface can be extended further in the $\rho_{o}/\rho_{c}\gg 1$ direction, where we find a maximum compactness value of $2M/R\approx 0.9706$. We have included a Mathematica notebook containing the pressure, mass, and redshift functions of both the inner and outer layers, as well as this 3D plot, so that any interested reader can visualize it from other angles.
By varying the values of $\rho_{o}$ and $R_{i}$, we can map out a surface in the three-dimensional parameter space $\{R_{i},\rho_{o},M\}$ that precisely represents the value of $M_{\infty}(R_{i},\rho_{o})$ for each pair $\left(R_{i},\rho_{o}\right)$. This surface is shown in Fig. 2. The vertical axis represents the maximum mass allowed for a given $R_{i}$ and $\rho_{o}$. All points on this surface lie between two constant $2M/R$ planes: one above Buchdahl’s limit ($92M/R=8/$ 9), which it intersects in the limits $\rho_{o}/\rho_{c}\to 1$ and $R_{i}/R\to 1$, and the other below the black hole limit ($12M/R=$ 1), which it neither intersects nor approaches asymptotically. Extending Fig. 2 in the direction of $\rho_{o}/\rho_{c}\gg 1$, we find the maximum value allowed for the compactness in the model, $C_{\text{max}}$, corresponding to the highest point in the 3D plot. The specific value is
@@
We have constructed a model in which the energy density grows outwards (albeit in a step-function-like fashion) while remaining always positive, thus breaking the second assumption from Buchdahl’s theorem. Despite this, we found it is not possible to arbitrarily approach the black hole limit $12M/R=$ 1 while keeping the energy density positive everywhere. Instead, the stars considered here display a maximum compactness $C_{\text{max}}$ beyond which the interior becomes singular. Figure 3 shows how this maximum compactness limit scales with the quotient between the outer and critical densities. The maximum appears around the value $\log\left(\rho_{o}/\rho_{c}\right)\approx 11.86$ for $R-R_{i}\approx 2.35\cdot 10^{-6}$. No additional maxima were found in our numerical explorations. We believe that this result is actually robust in the sense that it is independent of the step-function-like profile under consideration, its only model-dependent feature being the specific value of the maximum compactness bound, and not its overall existence.
Figure 3: Maximum compactness limit as a function of the quotient between the energy density of the outer layer and the critical density for stars with $\rho_{i}>0$. A single maximum compactness value of $2M/R\approx 0.9706$, shown in detail, is reached for $\log\left(\rho_{o}/\rho_{c}\right)\approx 11.86$ and $R-R_{i}\approx 2.35\cdot 10^{-6}$. No further maxima were found numerically for larger density values. For $\rho_{o}/\rho_{c}\to 1$, the compactness limit tends to the Buchdahl limit.
@@ III.2 Bilayered stars with vanishing core densities
A particularly illustrative case is the one in which the density of the inner core is zero. We treat it separately because its simplicity allows to find the compactness limit in an easier way. In this situation, the Misner-Sharp mass vanishes everywhere inside the core, and, for a fixed value of the density $\rho_{o}$ and the mass $M$, Eq. (27) enforces the radius $R_{i}$ to be precisely
@@
where we recall that the expression of $\rho_{c}$ in terms of $\{M,R\}$ is $\rho_{c}=3M/4\pi R^{3}$. This constraint also restricts pressure to be a positive, inwards-increasing monotonic function. This is because, starting from a positive, inwards-growing pressure, the numerator of the TOV equation (9) is only allowed to change sign if $m(r)<0$. Actually, we can explicitly integrate the TOV equation in the inner layer and find the following pressure profile
@@
where $\kappa$ is an arbitrary integration constant whose value is obtained from the matching to the outer layer, i.e.,
@@
In Subsection III.1 we noticed that the regular bilayered star with positive density that exhibits the largest possible compactness $C_{\text{max}}$ consisted of a narrow and highly dense shell surrounding an interior of (nearly) vanishing mass. In the present case where the inner core has exactly zero mass, the maximum compactness is attained by the solution (34) with $\kappa=0$, precisely the one that makes the pressure profile diverge exactly at $r=0$. To find this compactness bound, we calculate how $p_{i}$ scales in the $\rho_{o}/\rho_{c}\to\infty$ limit (the limit in which the outer shell is arbitrarily thin) for different values of the compactness.
@@
We have not been able to obtain an analytic expression but we have clearly seen numerically that $p_{i}$ always approaches a constant positive value, which we denote as $P_{i}^{~{}(\text{max})}$, as we vary $\rho_{o}$ while keeping the mass fixed for $\rho_{o}\gg\rho_{c}$. The particular value of $P_{i}^{~{}(\text{max})}$ is also found to increase with the mass, eventually diverging in the $2M/R\to 1$ limit, in agreement with the staticity requirement from Sec. II. From Eq. (35), we see that such divergent behaviour means that $\kappa\leq 0$ when $2M/R\to 1$. By calculating $P_{i}^{~{}(\text{max})}$ for a range of compactness values between the Buchdahl and black hole limits, we find that the compactness for which $\kappa=0$ agrees up to seven decimal places with the maximum compactness value (32) found for bilayered stars with positive density interiors.
In light of these results, we conjecture that given an outwards-increasing, monotonic density profile, there exists an upper bound to the compactness that isotropic stars can achieve, as long as we restrict ourselves to non-negative densities. In order to contrast this hypothesis, we have explored numerically different families of equations of state with similar characteristics that will be reported elsewhere, and in every case we find that there is a maximum available compactness which is actually smaller than the compactness of the toy model examined here. However, by no means we claim that the bilayered star profile is optimal, since, for instance, it is possible to think about this toy model as the limit of a parametric family of density profiles interpolating two constant density regions when the width of the interpolating region vanishes. Currently we lack any arguments to assert that these families—with the interpolation region keeping a finite width—yield less compact objects than the bilayered, step-function model considered here.
Finally, whereas a bound of the form $12m(r)/r<$ 1 was found in Baumgarte and Rendall (1993) assuming positive energy density and isotropy (see Mars et al. (1996) for an interesting extension to the anisotropic case, where only regularity at $r=0$ and the condition $\rho+p_{r}+2p_{t}\geq 0$ are assumed), our findings here suggest that a more stringent bound could be provided, although we leave the proof (or rebuttal) of this conjecture for future work.
@@ III.3 Bilayered stars with negative densities in the core
We have seen that, by considering a non-negative density profile that increases outward, the upper bound for the compactness exceeds the Buchdahl limit. However, there remains a gap between this new bound and the black hole limit. The next question that we pose is whether allowing the inner density $\rho_{i}$ to take negative values closes that gap, i.e., permits the existence of objects as compact as desired up to the black hole limit. Notice that this relaxation opens the door to violations of the weak (and hence the dominant) energy conditions.
In the case examined previously, we found that the solution which saturates the maximum compactness limit given in Eq. (32) looks (approximately) like a star with all its positive mass stored in a very narrow outer layer. The inner layer, on the contrary, occupies the whole bulk of the star and does not contribute to the mass at all. Such arrangement of the two layers is optimal in the sense that it maximizes the mass of the configuration for a fixed value of the outer density while varying $R_{i}$. A complementary interpretation is the following: the existence of a limit to the compactness implies that there is an optimal size for the outer layer. If this layer were made arbitrarily narrow, it would cause the central pressure to diverge. On the contrary, allowing for a negative density core enables the construction of stars for which the configuration that minimizes the central pressure does not require a very narrow and highly dense outer layer. Examples of these solutions are found by selecting the position of $R_{i}$ in such a way that the pressure remains almost constant inside the core. The compactness of solutions within this second family can be arbitrarily close to the black hole limit. The rest of this subsection is devoted to prove this statement in detail.
Assume, again, a bilayered model with an internal layer of negative density $\rho_{i}$. In order to proceed with the proof, we need to take a step back and consider for a moment the uniform-density, single-layer model ($R_{i}=0$). By inspection of Eqs. (23) and (24), we see that, for $\rho_{o}>\rho_{c}$, the Misner-Sharp mass $m(r)$ takes negative values in the interval $0\leq r<\left(1-\rho_{c}/\rho_{o}\right)^{1/3}R$. Extrapolated to $r=0$, this would point to the existence of a negative delta contribution to the density at the core $r=0$, meaning that the extrapolation of this configuration all the way to the centre of the star is not regular. Here, we only use this irregular configuration to generate the outer layer of the total geometry. In these configurations it is easy to see that the width of the negative-mass region grows with $\rho_{o}$. Now, by the TOV equation (9) we see that, at the surface $r=R$, pressure always increases inwards:
@@
Moreover, a negative Misner-Sharp mass can produce a maximum in the pressure at some $r=r_{\text{max}}$, which, by the Eq. (9), corresponds with the point when $m(r_{\text{max}})=-4\pi r_{\text{max}}^{3}p(r_{\text{max}})$. Values of $\rho_{o}$ bigger than a critical value $\rho_{\text{sep}}$ (whose specific value can be found numerically) ensure that it is possible to generate configurations with pressure profiles that are everywhere finite, in the same way that it is done in Arrechea et al. (2021): starting from $p(R)=0$ at the surface, then reaching a maximum value at $r=r_{\text{max}}$, to finally decrease until the centre $r=0$. In the $r\to 0$ limit, we obtain the following behaviour:
$p\simeq-\rho+k^{-2}\sqrt{\frac{r}{k}},\quad k>0.$ (37)
@@
which do not have real roots if $\rho_{i}<0$. On the other hand, the choice $R_{i}=R_{-}$ leads to an inner core that is characterized by a constant redshift function $\Phi(r)=\Phi_{i}$ and which satisfies the condition $03p+\rho_{i}=$ 0. Figures 4 and 5 show plots of these two solutions with constant-pressure inner cores. For any choice of $R_{i}$ in between those two values, namely $R_{-}<R_{i}<R_{+}$, we obtain regular solutions with $p^{\prime}>0$ inside the core. Regular solutions can also be found for $R_{i}\gtrsim R_{+}$ and for $R_{i}\lesssim R_{-}$ although we do not focus on them here since our aim is only to show that it is possible to find solutions as compact as desired.
Figure 4: Typical pressure and density profiles for stars displaying a constant-pressure inner layer such that $p+\rho_{i}=0$. The star shown in this Figure has $1.42M/R=0.99,~{}\rho_{o}=$ 1.4 and $R_{i}=R_{+}\approx 0.72$. In the limit where the outer layer is infinitesimally thin, these solutions reduce to the AdS star model (see Sec. IV below). It is always possible to find stars belonging to this family for any compactness $12M/R<$ 1.
Figure 5: Typical pressure and density profiles for stars displaying a constant-pressure inner layer such that $03p+\rho_{i}=$ 0. The star shown in this Figure has $1.42M/R=0.99,~{}\rho_{o}=$ 1.4 and $R_{i}=R_{-}\approx 0.53$. In the limit where the outer layer is infinitesimally thin, these solutions reduce to the Einstein static star model (see Sec. IV below). It is always possible to find stars belonging to this family for any compactness $12M/R<$ 1.
@@
As a summary, the main consequence of not constraining the sign of $\rho_{i}$ is that we allow to explore a broader region of the space of parameters $\{R_{i},\rho_{o},M\}$. This is depicted in Fig. 6 where, for some fixed mass value, only solutions inside the green region are allowed to have $\rho_{i}>0$. The green curve $R_{i}=\left(1-\frac{\rho_{c}}{\rho_{o}}\right)^{1/3}R$, where $\rho_{i}=0$, acts as the boundary of this region. Within the green portion of the diagram we find the maximum compactness bound (32). Once we move below said curve, thus allowing for solutions with $\rho_{i}<0$, it is always possible to find broad regions (shaded in blue) of perfectly regular stars, bounded by $R_{+}$ and $R_{-}$. The maximum compactness of these stars can be arbitrarily close to the black hole compactness.

### v2

@@ III.1 Bilayered stars with positive core densities
Figure 3: Three dimensional representation of the surface in the $\{R_{i},\rho_{o},M\}$ space corresponding to solutions with an infinite central pressure and $\rho_{i}>0$. Every point on this surface lies above the standard Buchdahl limit $92M/R=8/$ 9, which is reached in the $\rho_{o}/\rho_{c}\to 1,~{}R_{i}/R\to 1$ limits. Lighter colours denote larger values of $2M/R$, and the black contour lines correspond to the $2M/R=\{0.92,0.94,0.96\}$ planes. This surface can be extended further in the $\rho_{o}/\rho_{c}\gg 1$ direction where we find an asymptotic approach to the maximum compactness value of $2M/R\approx 0.9706$. We have included a Mathematica notebook containing the pressure, mass, and redshift functions of both the inner and outer layers, as well as this 3D plot, so that any interested reader can visualize it from other angles.
By varying the values of $\rho_{o}$ and $R_{i}$, we can map out a surface in the three-dimensional parameter space $\{R_{i},\rho_{o},M\}$ that precisely represents the value of $M_{\infty}(R_{i},\rho_{o})$ for each pair $\left(R_{i},\rho_{o}\right)$. This surface is shown in Fig. 3. The vertical axis represents the maximum mass allowed for a given $R_{i}$ and $\rho_{o}$. All points on this surface lie between two constant $2M/R$ planes: one above Buchdahl’s limit ($92M/R=8/$ 9), which it intersects in the limits $\rho_{o}/\rho_{c}\to 1$ and $R_{i}/R\to 1$, and the other below the black hole limit ($12M/R=$ 1), which it neither intersects nor approaches asymptotically. Extending Fig. 3 in the direction of $\rho_{o}/\rho_{c}\gg 1$, we find that $C(R)$ saturates at a maximum value which approximately is
@@
To the level of our numerical precision it agrees with Bondi bound (27). In fact, we find that there is no gap between this uppermost bound (40) and the Bondi bound. Figure 4 shows how this maximum compactness limit scales with the quotient between the outer and critical densities. Finally, in Figure 5 we have plotted three solutions belonging to this family with positive, albeit small, inner-layer densities.
Figure 4: Maximum compactness limit as a function of the quotient between the energy density of the outer layer and the critical density for stars with $\rho_{i}>0$. For $\rho_{o}/\rho_{c}\to 1$, the compactness limit tends to the Buchdahl limit whereas for $\rho_{o}/\rho_{c}\to\infty$ it approaches the Bondi limit.
Figure 5: $(u,v)$ diagram of bilayered stars with non-negative energy densities. We have represented the $v(u)$ curves from the inner layer in green and those from the outer layer in blue. As the compactness approaches the Bondi compactness, the bilayered model acquires characteristics very similar to Bondi’s model I, namely an inner layer that lies on the vertical axis and an outer layer that approaches the $P_{A}$-curve with $A=9$, which is the dashed orange line in the figure. Green curves closer to the vertical axis correspond to smaller inner-layer densities.
@@ III.2 Bilayered stars with vanishing core densities
A particularly illustrative case arises when the density of the inner core is zero, a feature also present in Model I proposed by Bondi. We treat it separately because its simplicity allows to find the compactness limit in an easier way. Moreover, we can provide numerical evidence for the absence of any finite gap between the maximum compactness attained by this bilayered model and Bondi’s bound. In the present situation, the Misner-Sharp mass vanishes everywhere inside the core and, for a fixed value of the density $\rho_{o}$ and the mass $M$, Eq. (35) enforces the radius $R_{i}$ to be precisely
@@
where we recall that the expression of $\rho_{c}$ in terms of $\{M,R\}$ is $\rho_{c}=3M/4\pi R^{3}$. This constraint also restricts the pressure to be a positive, inwards-increasing monotonic function. This is because, starting from a positive pressure that grows inward, the numerator of the TOV equation (9) is only allowed to change sign if $m(r)<0$. Actually, we can explicitly integrate the TOV equation in the inner layer and find the following pressure profile:
@@
where $\kappa$ is an arbitrary integration constant, corresponding to the inverse of the pressure at the origin, whose value is obtained from the matching with the outer layer, i.e.,
@@
In Subsection III.1 we noticed that the regular bilayered star with positive density that exhibits the largest possible compactness $C_{\text{max}}$ would be obtained in the limit of a thin and infinitely dense shell surrounding an interior of (nearly) vanishing mass. In the present case, where the inner core has exactly zero mass, the maximum compactness is attained by the solution (42) with $\kappa=0$, precisely the one that makes the pressure profile diverge exactly at $r=0$. To find this compactness bound, we calculate how $p_{i}$ scales in the $\rho_{o}/\rho_{c}\to\infty$ limit (the limit in which the outer shell is arbitrarily thin) for different values of the compactness.
@@
In our numerical explorations, we see that $p_{i}$ always approaches a constant positive value, which we denote as $P_{i}^{~{}(\text{max})}$, as we vary $\rho_{o}$ while keeping the mass fixed for $\rho_{o}\gg\rho_{c}$. By calculating $P_{i}^{~{}(\text{max})}$ for a range of compactness values between the Buchdahl and black hole limits, we find that, if the density of the outer layer is made very large, the value $\kappa=0$ is found very close to the Bondi limit, defined in Eq. (27). In particular, we find the scaling
$C(R)\approx C_{\rm{B}}-\frac{\alpha}{R^{2}\rho_{o}},$ (44)
where $\alpha$ is a dimensionless positive constant. The Bondi bound is only reached in the singular limit $\rho_{o}\to\infty$, which would correspond to putting all the mass of the star inside a thin shell—since, in virtue of (41), it implies that $R_{i}\to R$. Although the solution with $\kappa=0$ is, strictly speaking, singular, we can find solutions whose compactness approaches Bondi’s bound as much as desired through a physical model that is purely isotropic and can be appropriately matched with a Schwarzschild exterior.
Comparison with Bondi’s Model I.
In Figure 5, we have plotted two configurations in the $(u,v)$ variables belonging to our family of vanishing core densities. As the external crust becomes thinner, the compactness approaches Bondi’s bound more closely. The portion of the curve with a negative slope, where $v$ decreases as $u$ increases, corresponds to the thick shell region. Regardless of the shell thickness, these curves never coincide with the parabolic curves defined by a fixed parameter $A$. Instead, they consistently move downward, intersecting parabolas with different values of $A$, which is the crucial difference with respect to Bondi’s Model I.
Model I is described as a vanishing energy core with non-vanishing pressure connecting through a thin envelope to an external Schwarzschild geometry. The vanishing energy core is represented by the curve $u=0$ and $v\in[0,2-\epsilon]$, $\epsilon>0$, where
$\rho=0,\qquad p=\frac{p_{c}}{1+2\pi p_{c}r^{2}},\qquad 2\pi p_{c}r^{2}=\frac{v}{2-v},\qquad 0\leq r\leq\sqrt{\frac{2-\epsilon}{2\pi p_{c}\epsilon}},$ (45)
with $p_{c}$ being the value of the pressure at the centre $r=0$. The envelope is represented by $v=(3-\epsilon)(1-2u)^{1/2}-(1-u)$ and $u\in[0,u_{s}]$, with
$u_{s}=\left(3-\epsilon\right)(8-6\epsilon+\epsilon^{2})^{\frac{1}{2}}-\left(8-6\epsilon+\epsilon^{2}\right)=6\sqrt{2}-8-\order{\epsilon}.$ ) . (46)
As the envelope is strictly following one of the parabolic curves, it represents from the start an infinitely-thin limit. This is so because along such curve $r$ does not vary, and the energy density becomes infinite (cf. Eqs. (22)). By making $\epsilon$ arbitrarily small, the model can approach the Bondi compactness limit as closely as desired.
To better understand Bondi’s Model I, let us perform a detailed thin-shell analysis following Israel’s formalism: Take a spherical shell with radius $R$, such that its proper area is $4\pi R^{2}$. Let the spacetime be described by a external Schwarzschild patch that we denote with a $+$ sign and the internal patch that we denote with a -- sign. Generically, the metric reads:
$ds^{2}_{+}=-f_{+}(r_{+})dt_{+}^{2}+h_{+}(r_{+})dr_{+}^{2}+r_{+}^{2}d\Omega_{2}^{2},$ (47)
$ds^{2}_{-}=-f_{-}(r_{-})dt_{-}^{2}+h_{-}(r_{-})dr_{-}^{2}+r_{-}^{2}d\Omega_{2}^{2}.$ (48)
The external Schwarzschild patch corresponds to $f_{+}(r_{+})=1-2M/r_{+}$ and $h_{+}(r_{+})=1/f(r_{+})$. The internal geometry has a vanishing mass function and, as such, $h(r_{-})=1$ and the redshift function is given by
$f(r_{-})=(2\pi r_{-}^{2}+1/p_{c})^{2}.$ (49)
We locate the thin shell at $r_{+}=r_{-}=R$. The matching is worked out in detail in Appendix B, and we find that the shell displays a distributional stress-energy tensor with the following densities and pressures:
$\rho=\sigma\delta(r-R),$ (50)
$p_{r}=0,$ (51)
$p_{t}=\tilde{p}_{t}\delta(r-R),$ (52)
with $\sigma$ and $\tilde{p}_{t}$ given by
$\sigma$ $=\frac{1}{4\pi R}\left(1-\sqrt{1-\frac{2M}{R}}\right),$
$\tilde{p}_{t}$ $=\frac{1}{8\pi R}\left(\sqrt{1-\frac{2M}{R}}-1\right)+\frac{1}{16\pi R}\left(\frac{2M}{R}\frac{1}{\sqrt{1-\frac{2M}{R}}}-\frac{8\pi p_{c}R^{2}}{(1+2\pi p_{c}R^{2})}\right).$ (53)
Now, for a given value of the compactness, there is always a specific value of $p_{c}$ such that the tangential pressure vanishes, i.e., $\tilde{p}_{t}=0$, which corresponds to
$p_{c}=-\frac{R\left(\sqrt{1-\frac{2M}{R}}-1\right)+M}{2\pi R^{2}\left(R\left(3\sqrt{1-\frac{2M}{R}}-1\right)+M\right)}.$ (54)
In this scenario, the shell is held in place by the pressure from the matter content of the inner region, despite the shell itself being pressureless. This situation is peculiar because it involves an interior composed of unusual matter that exerts pressure without possessing any energy density. We discover that the compactness cannot be increased to the black hole limit, as there is a specific value at which the solution becomes singular for $r=0$, since the pressure there becomes infinite, i.e., $p_{c}\to\infty$. Increasing the compactness only moves this singularity to finite radii. By looking at Eq. (67), one can check that the threshold value of $p_{c}=+\infty$ corresponds precisely to a compactness
$C(R)=12\sqrt{2}-16,$ (55)
which is precisely Bondi’s bound $C(R)=C_{B}$. Thus, we have seen that Bondi’s Model I corresponds to the distributional limit of our models, with each providing further insights into the other. In fact, for any given compactness $C_{1}(R)$ lying between Buchdahl’s and Bondi’s limits (i.e., $C_{\rm Buchdahl}<C_{1}(R)<C_{\rm Bondi}$), there is always a bilayered model with that compactness, as well as vanishing density cores and sufficiently thin crusts which ensure that the integration of the TOV equation from the surface inwards leads to a completely regular configuration.
@@ III.3 Bilayered stars with negative densities in the core
We have demonstrated that an outward-increasing, non-negative density profile allows for compactness exceeding the Buchdahl limit, as established by Bondi, and we have constructed a toy model that nearly saturates this bound. Bondi further argues that permitting negative inner densities $\rho_{i}$ removes this restriction, allowing for configurations approaching the black hole compactness limit. In this Subsection, we provide an explicit example within the bilayered model, showing that for negative of $\rho_{i}$, objects can be constructed arbitrarily close to this limit. However, this relaxation also permits violations of the weak (and consequently the dominant) energy conditions.
In the case examined previously, we found that the solution which maximizes the compactness looks (approximately) like a star with all its positive mass stored in a very narrow outer layer. The inner layer, on the contrary, occupies the whole bulk of the star and does not contribute to the mass at all. Such arrangement of the two layers is optimal in the sense that it maximizes the mass of the configuration. Instead, allowing for a negative density core enables the construction of stars for which the configuration that minimizes the central pressure does not require a very narrow and highly dense outer layer. Examples of these solutions are found by selecting the position of $R_{i}$ in such a way that the pressure remains almost constant inside the core. The compactness of these solutions can be arbitrarily close to the black hole limit. The rest of this subsection is devoted to prove this statement in detail.
Assume a bilayered model with an internal layer of negative density $\rho_{i}$. In order to proceed with the proof, we need to take a step back and consider for a moment the uniform-density, single-layer model ($R_{i}=0$). By inspection of Eqs. (31) and (32), we see that, for $\rho_{o}>\rho_{c}$, the Misner-Sharp mass $m(r)$ takes negative values in the interval $0\leq r<\left(1-\rho_{c}/\rho_{o}\right)^{1/3}R$. Extrapolated to $r=0$, this would point to the existence of a negative delta contribution to the density at the core $r=0$, meaning that the extrapolation of this configuration all the way to the centre of the star is not regular. Here, we only use this irregular configuration to generate the outer layer of the total geometry. In these configurations it is easy to see that the width of the negative-mass region grows with $\rho_{o}$. Now, by the TOV equation (9) we see that, at the surface $r=R$, pressure always increases inwards:
@@
Moreover, a negative Misner-Sharp mass can produce a maximum in the pressure at some $r=r_{\text{max}}$, which, by the Eq. (9), corresponds with the point when $m(r_{\text{max}})=-4\pi r_{\text{max}}^{3}p(r_{\text{max}})$. Values of $\rho_{o}$ bigger than a critical value $\rho_{\text{sep}}$ (whose specific value can be found numerically) ensure that it is possible to generate configurations with pressure profiles that are everywhere finite, in the same way that it is done in Arrechea et al. (2021): starting from $p(R)=0$ at the surface, then reaching a maximum value at $r=r_{\text{max}}$, to finally decrease until the centre $r=0$. In the $r\to 0$ limit, we obtain the following behavior:
$p\simeq-\rho+k\sqrt{r},\quad k>0.$ (57)
@@
which do not have real roots if $\rho_{i}<0$. On the other hand, the choice $R_{i}=R_{-}$ leads to an inner core that is characterized by a constant redshift function $\Phi(r)=\Phi_{i}$ and which satisfies the condition $03p+\rho_{i}=$ 0. Figures 6 and 7 show plots of these two solutions with constant-pressure inner cores. For any choice of $R_{i}$ in between those two values, namely $R_{-}<R_{i}<R_{+}$, we obtain regular solutions with $p^{\prime}>0$ inside the core. Regular solutions can also be found for $R_{i}\gtrsim R_{+}$ and for $R_{i}\lesssim R_{-}$ although we do not focus on them here since our aim is only to show that it is possible to find solutions as compact as desired.
Figure 6: Typical pressure and density profiles for stars displaying a constant-pressure inner layer such that $p+\rho_{i}=0$. The star shown in this Figure has $1.42M/R=0.99,~{}R^{2}\rho_{o}=$ 1.4 and $R_{i}=R_{+}\approx 0.72R$. In the limit where the outer layer is infinitesimally thin, these solutions reduce to the AdS star model (see Sec. IV below). It is always possible to find stars belonging to this family for any compactness $12M/R<$ 1.
Figure 7: Typical pressure and density profiles for stars displaying a constant-pressure inner layer such that $03p+\rho_{i}=$ 0. The star shown in this Figure has $1.42M/R=0.99,~{}R^{2}\rho_{o}=$ 1.4 and $R_{i}=R_{-}\approx 0.53R$. In the limit where the outer layer is infinitesimally thin, these solutions reduce to the Einstein static star model (see Sec. IV below). It is always possible to find stars belonging to this family for any compactness $12M/R<$ 1.
@@
As a summary, the main consequence of not constraining the sign of $\rho_{i}$ is that we allow to explore a broader region of the space of parameters $\{R_{i},\rho_{o},M\}$. This is depicted in Fig. 8 where, for some fixed mass value, only solutions inside the green region are allowed to have $\rho_{i}>0$. The green curve $R_{i}=\left(1-\frac{\rho_{c}}{\rho_{o}}\right)^{1/3}R$, where $\rho_{i}=0$, acts as the boundary of this region. Within the green portion of the diagram we find the maximum compactness bound (40). Once we move below said curve, thus allowing for solutions with $\rho_{i}<0$, it is always possible to find broad regions (shaded in blue) of perfectly regular stars, bounded by $R_{+}$ and $R_{-}$. The maximum compactness of these specific stars can be arbitrarily close to the black hole compactness.
These solutions have a complementary interpretation in terms of the $(u,v)$ variables used in Subsection II.2. The $v(u)$ curve describing the constant-density core solutions is given by
$v=-\frac{u\left(v_{0}+3\sqrt{3-6u}\right)}{v_{0}+\sqrt{3-6u}},$ (61)
where $v_{0}$ is an integration constant related to the central pressure via
$v_{0}=-\frac{3\sqrt{3}\left(\rho_{i}+p(0)\right)}{\rho_{i}+3p(0)}$ (62)
The constant-pressure curves $\rho+p_{i}=0$ and $\rho+3p_{i}=0$ are just the straight lines $v=-3u$ and $v=-u$ in the $(u,v)$ diagram from Fig. 1. These are obtained by evaluating (61) in the limits $v_{0}\to 0$ and $v_{0}\to\pm\infty$, respectively. Since these $v$ lines always intersect the $P_{A}$ parabolas (26) at some $u<0$ for any $A>1$, it is possible to match them to an outer constant-density layer so that the compactness approaches the black hole compactness $u=1/2$ $(A\to\infty)$ as much as desired. Fig. 9 shows the $v(u)$ curves for the two models with constant-pressure inner layers that we presented above.
Figure 9: $(u,v)$ diagram of bilayered stars with negative energy densities. The $v(u)$ curves from the inner layer appear in green and those from the outer layer in blue. We have plotted two example solutions satisfying $\rho_{i}+p=0$ and $\rho_{i}+3p=0$ and which surpass Bondi’s limit. The compactness of these solutions can be arbitrarily close to the black hole limit.